Five uncharged capacitors with equal capacitances are combined in parallel. The combination is connected to a 5.25 V battery, which charges the capacitors. The charging process involves 0.000243 C of charge moving through the battery. Find the capacitance of each capacitor. capacitance:
Answer #1

Since the capacitors are in parallel, the Total capacitance = C1 + C2 + C3 + C4 + C5. Since the capacitance of each capacitor is equal. Total Capacitance= 5C.

Now Q=CV, C=Q/V

hence 5C= 0.000243/5.25

Hence capacitance of each capacitor = C = 9.2 8 10^(-6) farad